OpenStudy (anonymous):

Solve for x: 2/x = 6/2

OpenStudy (anonymous):

cross multiply

jimthompson5910 (jim_thompson5910):

2/x = 6/2 2*2 = x*6 .. cross multiply 4 = 6x ... ... ... x = ???

OpenStudy (anonymous):

it would come out to be 1.5 or 1 1/2

jimthompson5910 (jim_thompson5910):

that's incorrect

OpenStudy (anonymous):

the answers i have to select is 2/3. 1 1/2 or 6

jimthompson5910 (jim_thompson5910):

Nurali you have things flipped

OpenStudy (anonymous):

It would be 2/3 then?

jimthompson5910 (jim_thompson5910):

x = 2/3 is correct

OpenStudy (anonymous):

Okay thank you. Im really bad at these,

OpenStudy (nurali):

2/x = 6/2 6x=4 x=4/6 x=2/3